If you had an empty cylinder 10 inches in diameter and 20 inches long and applied 10 psi of air pressure to the inside, what would the force be on one of the ends of the cylinder?
Answer by Alan3354(69443)

Force = Pressure * Area
------------
Area of 10" = pi*r^2 = 25pi sq inches
----------------
F = 10*25pi = 250pi pounds =~ 785.4 pounds on both ends
----------